Suppose the company is expected to pay a dividend of $8.97. next year That is, D1 is $8.97. What is the amount of dividend in Year 20? That is, what is D20? Assume that the dividends are expected to grow by 15% each year.

Enter your answer rounded off to two decimal points. Do not enter $ or comma in the answer box.
Answer #1

First of all lets find Dividend of year 0
Dividend of year 1 = Dividend of year 0(1+growth rate)
8.97 = Dividend of year 0(1+15%)
8.97 = Dividend of year 0(1+0.15)
8.97 = Dividend of year 0(1.15)
Dividend of year 0 = 7.8 $
Now we can use formula of Future value
we have Present value = PV = 7.8 $ , n = no of years = 20 , r =15%
FV =PV(1+r)^n
=7.8(1+0.15)^20
=7.8(1.15)^20
=7.8(16.3665)
=127.66 $